# From -gram to Attention: How Model Architectures Learn and Propagate Bias in Language Modeling

### v1

@@ Abstract
From -gram to Attention: How Model Architectures Learn and Propagate Bias in Language Modeling
Mohsinul Kabir, Tasfia Tahsin, Sophia Ananiadou Department of Computer Science, National Center for Text Mining, The University of Manchester Department of Computer Scinece and Engineering, Islamic University of Technology {mdmohsinul.kabir, sophia.ananiadou}@manchester.ac.uk, tasfiatahsin@iut-dhaka.edu
@@ 1 Introduction
The vast scale and often unverified quality of training data (heterogeneous data) create substantial pathways for social biases to manifest in language models, whether in relatively simple word embeddings (Bolukbasi etΒ al., 2016; Basta etΒ al., 2019) or in far more complex large language models (Abid etΒ al., 2021; Abrar etΒ al., 2025). Such harmful biases pose significant challenges to the deployment of large-scale AI systems in diverse sociocultural contexts Kabir etΒ al. (2025). This growing awareness has spurred a considerable body of NLP research focused on measuring and mitigating bias. However, the relationship between bias and language models remains complex, and the mechanisms underlying biased decision-making in downstream tasks are not yet fully understood (Kruspe, 2024). Contemporary attempts at bias mitigation have often been criticized as βputting lipstick on a pigβ, meaning that they merely conceal bias rather than truly eliminating it (Gonen and Goldberg, 2019). This criticism largely stems from the fact that current approaches to bias identification and reduction tend to operate as black-box methods, offering limited transparency. Most studies attribute bias primarily to web data used for training language models (Baeza-Yates, 2016), while some studies have sought to connect the model architecture with observed biases (Liu etΒ al., 2024) . However, recent research often overlooks the complex interplay between training data and model architecture in shaping model biases (Vig etΒ al., 2020). Even the time of creation of the training corpora can significantly affect the biases induced in language models, a factor that remains largely unexplored at the experimental level (Navigli etΒ al., 2023). So, before we can effectively mitigate biases in language models, an essential first step is to βunderstandβ the bias.
@@
Modern language models are fundamentally autoregressive generators trained for language modeling tasks (Jurafsky and Martin, ). To understand how bias propagates across different model architectures, it is important to analyze this phenomenon through the lens of language modeling. Transformers, which form the foundation of many autoregressive generation models (Vaswani etΒ al., 2017), have been the subject of extensive research on interpretability (Borenstein etΒ al., 2024; Nowak etΒ al., 2024). In this context, comparative behavioral analysis frameworks that juxtapose transformers with -gram models have emerged as valuable tools to understand transformer behavior (Voita etΒ al., 2024; Svete etΒ al., 2024; Svete and Cotterell, 2024). Comparative behavioral analysis frameworks assess model behavior either by comparing performance under the same controlled conditions (Isabelle etΒ al., 2017; Naik etΒ al., 2018) or by visualizing important input features using saliency methods (Murdoch etΒ al., 2018). In this work, we investigate how comparative behavioral analysis supports the interplay between training data and model architecture and influences the propagation of bias during language modeling.
@@
Effects of architectural parameters: We empirically demonstrate how architectural design choices influence bias propagationβspecifically, context window size and smoothing techniques in -gram models versus layer depth, attention heads, and attention types in transformers. Our results indicate that transformers are more robust to contextual bias than -gram models.
@@
Temporal influence of training data: We examine how the temporal provenance of training data affects bias propagation in language models, showing that Wikipedia dumps from different years lead to distinct bias dynamics in the resulting models.
@@
Categorial bias preference: We investigate whether certain categories of social bias are more strongly amplified in specific model architectures. Our findings confirm differential amplification across bias types, underscoring the need for targeted mitigation strategies.
Addressing these aspects brings us closer to a practical understanding of bias propagation in language modeling, illuminating how the theoretical properties of neural architectures translate into real-world bias compared to -gram statistics. Our approach using comparative behavioral framework can be generalised to other architectures that use language modeling tasks, providing an interpretable tool for bias propagation and consequently paving the way for more effective bias mitigation in language models.
A comprehensive background review of contemporary research on -grams versus transformers, bias and interpretability is presented in AppendixΒ A.
@@ 2 Bias Propagation in LMs
At a high level, our work compares how bias emerges in statistical -gram LMs versus neural transformers during language modeling. We train both types of model with varied architectural parameters on Wikipedia data from different time periods, then systematically evaluate their biases using the CrowS-Pairs dataset (Nangia etΒ al., 2020).
@@ 2.1 Experimental Setup
Our methodology comprises three key components: (1) training data preparation using three Wikipedia data dumps (2018, 2020, 2024) with incremental injections of stereotypical examples (0%, and mixtures); (2) model training across architectural variants, including -gram, -gram, and -gram models with Kneser-Ney/Laplace/Add- smoothing for -grams, and transformer configurations varying in depth ( layers), attention heads (), and attention types (soft/sparse). These parameter choices are motivated by the experiments conducted by Svete etΒ al. (2024); and (3) rigorous bias quantification using the CrowS-Pairs dataset for stereotypical versus anti-stereotypical completions, along with measuring relative bias preferences among categories (gender, religion, race-color, etc.) .
@@ 2.2 Models
-gram Models. We train -gram language models with varying context windows () to examine how local context length influences bias propagation. These count-based models estimate next-token probabilities through maximum likelihood estimation of -gram frequencies in the training corpus (Jurafsky and Martin, ). Formally, for each sentence in the dataset, we construct the sequence and for each position , increase the count for the -gram and its context . During the evaluation, we apply three smoothing techniques: (1) Laplace (add-one) smoothing, (2) add- smoothing, and (3) modified Kneser-Ney interpolation. Each method implements distinct probability redistribution strategies for unseen -grams while regularizing observed counts (Ney etΒ al., 1994). For each Wikipedia dump (2018, 2020, 2024), we train three data variants: (i) the original unmodified corpus (0% bias), (ii) a 33% blend with synthetic stereotypical examples, and (iii) a fully augmented (100% blend) version.
Transformer Models. We analyze how architectural choices affect bias propagation in transformers by systematically varying two key parameters: (1) model depth (number of layers, ) and (2) attention head count (). Moreover, we examine how attention mechanisms influence bias dynamics by comparing softmax with sparsemax attention variants (Vaswani etΒ al., 2017). Each transformer model is trained using a standard autoregressive objective, optimizing the model parameters to maximize the likelihood of the next token given the preceding context, following the standard autoregressive factorization: where denotes the -th token and the model parameters. Mirroring our -gram experiments, each transformer configuration is trained on all Wikipedia dumps with identical bias injection levels (0%, 33%, and 100% synthetic data). The summary of the trained -gram and transformer models along with their parametric variations is presented in Table 1. Further details regarding our model training process are provided in Appendix B.
@@ 2.3 Data for LM Training
We choose data from three English Wikipedia data dumps to train our models: English Wikipedia Dump from August 2018, October 2020 and April 2024. Wikipedia data dumps are publicly available and serve a variety of purposes, including research, offline analysis, archiving, etc. Since Wikipedia constitutes a significant portion of LLM training corpora (Cheng etΒ al., 2024), analyzing these dumps across different timeframes allows us to study the temporal drift in bias and its potential impact on language models. We do not restrict our selection to any specific article categories; instead, our dataset encompasses content from the full range of Wikipedia articles to capture a diverse spectrum topics.
@@
While Wikipedia data dumps primarily contain factual content, large language models (LLMs) are typically trained on a much broader range of web data, which can include stereotypical or biased languageΒ (Acerbi and Stubbersfield, 2023). To better simulate real-world LLM training conditions and analyze bias, we supplement Wikipedia data with synthetically generated stereotypical sentences. Using gpt-4o, we generate samples covering ten bias categories, employing the expert prompting technique (Xu etΒ al., 2023) as illustrated in Figure 6. Details about the full training data are presented in Table 2. Although the proportion of synthetic bias data is very small compared to the Wikipedia dump, we argue that this reflects real-world scenarios, where subtle biases exist within vast and heterogeneous training corpora of language models (Guo etΒ al., 2024).
@@
| Types of Stereotypes | β | β | β | 10 |
Further details on the training data and the generated stereotypical instances are provided in Appendix 2.
@@ 2.4 Bias Evaluation Framework
Once the models are trained, we evaluate bias using the CrowS-Pairs dataset (Nangia etΒ al., 2020). CrowS-Pairs is a widely used benchmark for assessing social biases in language models (Bommasani etΒ al., 2023), and has also been extended to other languages111https://gitlab.inria.fr/corpus4ethics/multilingualcrowspairs. The dataset contains 1,508 samples spanning nine bias categories: race, gender/gender identity, sexual orientation, religion, age, nationality, disability, physical appearance, and socioeconomic status. Each sample in CrowS-Pairs consists of a pair of sentences: one expressing a stereotype () and the other expressing an anti-stereotype () or less stereotypical view about one of the nine bias categories. As our goal is to evaluate bias in the language modeling task, we employ an adapted but aligned version of the bias evaluation metric originally proposed by Nangia etΒ al. (2020).
@@
A model that does not incorporate stereotypes regarding the bias categories in the dataset, therefore, shows neutrality to biases, should achieve the ideal score of . A bias score significantly higher than indicates that the model has a tendency to favor stereotypical views over non-stereotypical ones, and vice versa for scores significantly below .
For each of the bias categories described in CrowS-Pairs, we track the percentage of examples in which the model prefers the stereotypical sentence, providing a detailed breakdown of bias-type preferences. We also perform appropriate statistical tests for each of our experiments to determine whether the differences in the model responses are significant.
@@ 3.1 Effect of Architectural Parameters
Tables 3 and 5 present the mean bias scores for all training data and the corresponding standard deviations for all model configurations. The -gram and transformer language models demonstrate distinct behaviors as their architectural parameters are varied. While -gram models tend to exhibit a more linear and predictable relationship between parameter changes and bias scores, transformers exhibit complex, non-linear patterns, reflecting the manifestation of their sophisticated architectural dynamics.
@@
Modified Kneser-Ney smoothing stands out for its neutral response to changes in -gram order, consistently maintaining bias scores very close to the ideal value of with minimal standard deviation across all -gram configurations. This finding aligns with previous research, which highlights the superior performance of Kneser-Ney smoothing to model rare and unseen events by leveraging multiple discount parameters and interpolating lower-order probabilities, resulting in lower perplexity compared to other smoothing methods (DobΓ³, 2018; James, 2000). Our findings further support its effectiveness, demonstrating that Kneser-Ney mitigates bias more efficiently than alternative techniques. FigureΒ 7 shows a staircase pattern of decrease in bias scores with higher -gram order for Laplace and add- smoothing, while Kneser-Ney smoothing consistently achieves near-ideal bias scores.
We compute Spearmanβs to assess the relationship between -gram order and bias score for each smoothing method. As shown in TableΒ 4, Laplace and add- show an almost perfect, highly significant negative correlation, indicating that bias scores consistently decrease as -gram order increases. For Kneser-Ney, this trend is weaker and only significant when bias is injected (wiki+ bias and wiki+ bias).
@@
Transformer models demonstrate exceptional resilience to bias across all examined hyperparameters. As shown in Table 5, both sparse and soft attention transformer architectures consistently yield bias scores near the ideal value of , regardless of the configuration. Figures 8 and 9 show that variations in the number of layers (, , ) and attention heads (, , ) do not result in systematic changes in bias scores. The violin plots presented in Figure 1 further illustrate the robustness of transformer models, as their bias scores remain tightly clustered across all heads and layers. This stands in stark contrast to the -gram models, which display greater susceptibility to bias, particularly at higher -gram orders, highlighting the superior bias stability of transformer-based approaches.
@@
The Spearmanβs (and corresponding -values) between bias score and both the number of layers and attention heads for each attention type are reported in TableΒ 6. All correlation coefficients are near zero and not statistically significant, indicating no meaningful association between these architectural parameters and bias scores for either attention mechanism.
@@
A close examination of TableΒ 5 reveals a subtle but consistent dominance of sparse attention over soft attention in terms of bias mitigation. To quantitatively assess which attention mechanism yields bias scores closer to the ideal value of , we compute the mean absolute deviation for both sparse and soft attention across all experimental settings using the following equation:
where denotes the bias score for each configuration and is the total number of configurations considered. Our results indicate that sparse attention achieves a mean deviation of , while soft attention yields . This finding demonstrates that, on average, the bias scores produced by sparse attention are nearly twice as close to perfect neutrality () as those generated by soft attention. This advantage is further illustrated in FigureΒ 1, where the KDE plots for sparse attention consistently cluster around the ideal bias score across all Wikipedia dumps, in contrast to the broader and more variable distributions observed for soft attention.
@@ 3.2 Temporal Influence of Data
We hypothesize that training data from different time frames induce bias differently in models, regardless of the underlying architecture, as also suggested by Navigli etΒ al. (2023). However, there is a lack of systematic evaluation of this effect. To address this, we analyze three Wikipedia dumps from 2018, 2020, and 2024 to investigate how the temporal characteristics of training data influence bias propagation. Specifically, we aim to determine whether data from different periods lead to varying levels of bias in the resulting models.
We conduct paired -tests between bias scores for all year pairs (2018, 2020, 2024) for both -gram and transformer models. We perform these analyses across three bias injection levels (0%, 33%, and 100%) to disentangle the effect of explicit bias injection from the temporal effect of the data itself. As shown in TableΒ 7, the 2018 Wikipedia dump inherently exhibits an anti-stereotype bias, as evidenced by mean bias scores for all models being well below the ideal unbiased score of . Focusing on the wiki_only rows, we observe a marked increase in bias scores for 2020 across all models, with transformer models (both soft and sparse attention) showing particularly pronounced increases, approaching the ideal score of . All of these increases are statistically significant. When a moderate amount of bias () is injected into the training data, the models display a similar pattern: a significant increase in bias scores from 2018 to 2020. In contrast, a significant decrease in bias scores emerges between 2020 and 2024. However, we do not observe a significant change in bias scores between 2018 and 2024, or when the Wikipedia dump is supplemented with synthetic bias. These findings are visually illustrated in FigureΒ 10. We conduct further experiments on raw training data (excluding model effects) that reveal significant bias variations between pre- and post-COVID corpora, as detailed in AppendixΒ C.
Over time, words acquire new senses (e.g., mouse, tweet), and the predominant meanings of some words shift considerably. This continual evolution of language, combined with the choice of training corpora, introduces a significant subtype of selection bias in language modelsΒ (Navigli etΒ al., 2023). Our findings underscore the critical role of temporal influence in bias propagation, highlighting the importance of considering temporal factors when training and evaluating language models.

### v2

@@ Abstract
From -gram to Attention: How Model Architectures Learn and
Propagate Bias in Language Modeling
Mohsinul Kabir, Tasfia Tahsin, Sophia Ananiadou Department of Computer Science, National Center for Text Mining, The University of Manchester Department of Computer Science and Engineering, Islamic University of Technology {mdmohsinul.kabir, sophia.ananiadou}@manchester.ac.uk, tasfiatahsin@iut-dhaka.edu
@@ 1 Introduction
The vast scale and often unverified quality of training data (heterogeneous data) create substantial pathways for social biases to manifest in language models, whether in relatively simple word embeddings (Bolukbasi et al., 2016; Basta et al., 2019) or in far more complex large language models (Abid et al., 2021; Abrar et al., 2025). Such harmful biases pose significant challenges to the deployment of large-scale AI systems in diverse sociocultural contexts Kabir et al. (2025). This growing awareness has spurred a considerable body of NLP research focused on measuring and mitigating bias. However, the relationship between bias and language models remains complex, and the mechanisms underlying biased decision-making in downstream tasks are not yet fully understood (Kruspe, 2024). Contemporary attempts at bias mitigation have often been criticized as ‘putting lipstick on a pig’, meaning that they merely conceal bias rather than truly eliminating it (Gonen and Goldberg, 2019). This criticism largely stems from the fact that current approaches to bias identification and reduction tend to operate as black-box methods, offering limited transparency. Most studies attribute bias primarily to web data used for training language models (Baeza-Yates, 2016), while some studies have sought to connect the model architecture with observed biases (Liu et al., 2024) . However, recent research often overlooks the complex interplay between training data and model architecture in shaping model biases (Vig et al., 2020). Even the time of creation of the training corpora can significantly affect the biases induced in language models, a factor that remains largely unexplored at the experimental level (Navigli et al., 2023). So, before we can effectively mitigate biases in language models, an essential first step is to ‘understand’ the bias.
@@
Modern language models are fundamentally autoregressive generators trained for language modeling tasks (Jurafsky and Martin, ). To understand how bias propagates across different model architectures, it is important to analyze this phenomenon through the lens of language modeling. Transformers, which form the foundation of autoregressive generation models (Vaswani et al., 2017), have been the subject of extensive research on interpretability (Borenstein et al., 2024; Nowak et al., 2024). In this context, comparative behavioral analysis frameworks that juxtapose transformers with -gram models have emerged as valuable tools to understand transformer behavior (Voita et al., 2024; Svete et al., 2024; Svete and Cotterell, 2024). These frameworks assess model behavior either by comparing performance under the same controlled conditions (Isabelle et al., 2017; Naik et al., 2018) or by visualizing important input features using saliency methods (Murdoch et al., 2018). In this work, we investigate how the comparative behavioral analysis framework illuminates the interplay between training data and model architecture and influences the propagation of bias during language modeling.
@@
Effects of architectural parameters: We empirically demonstrate how architectural design choices influence bias propagation, specifically, context window size and smoothing techniques in -gram models versus layer depth, attention heads, and attention types in transformers. Our results indicate that transformers are more robust to contextual bias than -gram models.
@@
Temporal influence of training data: We examine how the origin time of the training data affects bias propagation in language models, showing that Wikipedia dumps from different years lead to distinct bias dynamics in the resulting models.
@@
Categorical bias preference: We investigate whether certain categories of social bias are more strongly amplified in specific model architectures. Our findings confirm differential amplification across bias types, underscoring the need for targeted mitigation strategies.
Addressing these aspects brings us closer to a practical understanding of bias propagation in language modeling, illuminating how the theoretical properties of neural architectures translate into real-world bias compared to -gram statistics. Our approach using comparative behavioral framework can be generalized to other architectures that use language modeling tasks, providing an interpretable tool for bias propagation and consequently paving the way for more effective bias mitigation in language models.
A comprehensive background review of contemporary research on -grams versus transformers, bias and interpretability is presented in Appendix A.
@@ 2 Bias Propagation in LMs
At a high level, our work compares how bias emerges in statistical -gram LMs versus neural transformers during language modeling. We train both types of models with varied architectural parameters on Wikipedia data from different time periods and then systematically evaluate their biases using the CrowS-Pairs dataset (Nangia et al., 2020).
@@ 2.1 Experimental Setup
Our methodology comprises three key components: (1) training data preparation using three Wikipedia data dumps (2018, 2020, 2024) with incremental injections of stereotypical examples (0%, and mixtures); (2) model training across architectural variants, including -gram, -gram, and -gram models with Kneser-Ney/Laplace/Add- smoothing for -grams, and transformer configurations varying in depth ( layers), attention heads (), and attention types (soft/sparse). These parameter choices are motivated by the experiments conducted by Svete et al. (2024); and (3) rigorous bias quantification using the CrowS-Pairs dataset for stereotypical versus anti-stereotypical completions, along with measuring relative bias preferences among categories (gender, religion, race-color, etc.) .
@@ 2.2 Models
-gram Models. We train -gram language models with varying context windows () to examine how local context length influences bias propagation. These count-based models estimate next-token probabilities through maximum likelihood estimation of -gram frequencies in the training corpus (Jurafsky and Martin, ). Formally, for each sentence in the dataset, we construct the sequence and for each position , increase the count for the -gram and its context . During the evaluation, we apply three smoothing techniques: (1) Laplace (add-one) smoothing, (2) add- smoothing, and (3) modified Kneser-Ney interpolation. Each method implements distinct probability redistribution strategies for unseen -grams while regularizing observed counts (Ney et al., 1994). For each Wikipedia dump (2018, 2020, 2024), we train three data variants: (i) the original unmodified corpus (0% bias), (ii) a % blend with synthetic stereotypical examples, and (iii) a fully augmented (% blend) version.
Transformer Models. We analyze how architectural choices affect bias propagation in transformers by systematically varying two key parameters: (1) model depth (number of layers, ) and (2) attention head count (). Moreover, we examine how attention mechanisms influence bias dynamics by comparing softmax with sparsemax attention variants (Vaswani et al., 2017).
Mirroring our -gram experiments, each transformer configuration is trained on all Wikipedia dumps with identical bias injection levels (0%, 33%, and 100% synthetic data). The summary of the trained -gram and transformer models along with their parametric variations is presented in Table 1. Further details regarding our model training process are provided in Appendix B.
@@ 2.3 Data for LM Training
We choose data from three English Wikipedia data dumps to train our models: English Wikipedia dump from August 2018, October 2020 and April 2024. Wikipedia data dumps are publicly available and serve a variety of purposes, including research, offline analysis, archiving, etc. Since Wikipedia constitutes a significant portion of LLM training corpora (Cheng et al., 2024), analyzing these dumps across different timeframes allows us to study the temporal drift in bias and its potential impact on language models. We do not restrict our selection to any specific article categories; instead, our dataset encompasses content from the full range of Wikipedia articles to capture a diverse spectrum of topics.
@@
While Wikipedia data dumps primarily contain factual content, large language models (LLMs) are typically trained on a much broader range of web data, which can include stereotypical or biased language (Acerbi and Stubbersfield, 2023). To better simulate real-world LLM training conditions and analyze bias, we supplement Wikipedia data with synthetically generated stereotypical sentences. Using gpt-4o, we generate samples covering ten bias categories, employing the expert prompting technique (Xu et al., 2023) as illustrated in Figure 8. Details about the full training data are presented in Table 2. Although the proportion of synthetic bias data is very small compared to the Wikipedia dump, we argue that this reflects real-world scenarios, where subtle biases exist within vast and heterogeneous training corpora of language models (Guo et al., 2024). Further details on the training data and the generated stereotypical instances are provided in Appendix 2.
@@
| Types of Stereotypes | – | – | – | 10 |
@@ 2.4 Bias Evaluation Framework
Once the models are trained, we evaluate bias using the CrowS-Pairs dataset (Nangia et al., 2020). CrowS-Pairs is a widely used benchmark for assessing social biases in language models (Bommasani et al., 2023), and has also been extended to other languages111https://gitlab.inria.fr/corpus4ethics/multilingualcrowspairs. The dataset contains 1,508 samples spanning nine bias categories: race, gender/gender identity, sexual orientation, religion, age, nationality, disability, physical appearance, and socioeconomic status. Each sample in CrowS-Pairs consists of a pair of sentences: one expressing a stereotype () and the other expressing an anti-stereotype () or less stereotypical view about one of the nine bias categories. As our goal is to evaluate bias in the language modeling task, we employ an adapted but aligned version of the bias evaluation metric originally proposed by Nangia et al. (2020).
@@
A model that does not incorporate stereotypes from the various bias categories in the dataset, therefore, shows neutrality to biases, should achieve the ideal score of . A bias score significantly higher than indicates that the model has a tendency to favor stereotypical views over non-stereotypical ones, and vice versa for scores significantly lower than .
For each of the bias categories described in CrowS-Pairs, we track the percentage of examples in which the model prefers the stereotypical sentence, providing a detailed breakdown of bias-type preferences. We also perform appropriate statistical tests for each of the experiments to determine whether the differences in the model responses are significant.
@@ 3.1 Effect of Architectural Parameters
Modified Kneser-Ney smoothing stands out for its neutral response to changes in -gram order, consistently maintaining bias scores very close to the ideal value of with minimal standard deviation across all -gram configurations. This finding aligns with previous research, which highlights the superior performance of Kneser-Ney smoothing to model rare and unseen events by leveraging multiple discount parameters and interpolating lower-order probabilities, resulting in lower perplexity compared to other smoothing methods (Dobó, 2018; James, 2000). Our findings further support its effectiveness, demonstrating that Kneser-Ney mitigates bias more efficiently than alternative techniques. Figure 9 shows a staircase pattern of decrease in bias scores with higher -gram order for Laplace and add- smoothing, while Kneser-Ney smoothing consistently achieves near-ideal bias scores.
We compute Spearman’s to assess the relationship between -gram order and bias score for each smoothing method. As shown in Table 4, Laplace and add- show an almost perfect, highly significant negative correlation, indicating that bias scores consistently decrease as -gram order increases. For Kneser-Ney, this trend is weaker and only significant when bias is injected (wiki+ bias and wiki+ bias).
@@
Transformer models demonstrate exceptional resilience to bias across all examined hyperparameters. As shown in Table 5, both sparse and soft attention transformer architectures consistently yield bias scores near the ideal value of , regardless of the configuration. Figures 10 and 11 show that variations in the number of layers (, , ) and attention heads (, , ) do not result in systematic changes in bias scores. The violin plots presented in Figure 1 further illustrate the robustness of transformer models, as their bias scores remain tightly clustered across all heads and layers. This stands in stark contrast to the -gram models, which display greater susceptibility to bias, particularly at higher -gram orders, highlighting the superior bias stability of transformer-based approaches.
@@
The Spearman’s (and corresponding ‑values) between bias score and both the number of layers and attention heads for each attention type are reported in Table 6. All correlation coefficients are near zero and not statistically significant, indicating no meaningful association between these architectural parameters and bias scores for either attention mechanism.
@@
A close examination of Table 5 reveals a subtle but consistent dominance of sparse attention over soft attention in terms of bias mitigation. To quantitatively assess which attention mechanism yields bias scores closer to the ideal value of , we compute the mean absolute deviation for both sparse and soft attention across all experimental settings using the following equation:
where denotes the bias score for each configuration and is the total number of configurations considered. Our results indicate that sparse attention achieves a mean deviation of , while soft attention yields . This finding demonstrates that, on average, the bias scores produced by sparse attention are nearly twice as close to perfect neutrality () as those generated by soft attention. This advantage is further illustrated in Figure 1, where the KDE plots for sparse attention consistently cluster around the ideal bias score across all Wikipedia dumps, in contrast to the broader and more variable distributions observed for soft attention.
@@ 3.2 Temporal Influence of Data
We hypothesize that training data from different time frames induce bias differently in models, regardless of the underlying architecture, as also suggested by Navigli et al. (2023). However, there is a lack of systematic evaluation of this effect. To address this, we analyze three Wikipedia dumps from 2018, 2020, and 2024 to investigate how the temporal characteristics of training data influence bias propagation. Specifically, we aim to determine whether data from different periods lead to varying levels of bias in the resulting models.
We conduct paired -tests between bias scores for all year pairs (2018, 2020, 2024) for both -gram and transformer models. We perform these analyses across three bias injection levels (0%, 33%, and 100%) to disentangle the effect of explicit bias injection from the temporal effect of the data itself. As shown in Table 7, the 2018 Wikipedia dump inherently exhibits an anti-stereotype bias, as evidenced by mean bias scores for all models being well below the ideal unbiased score of . Focusing on the wiki_only rows, we observe a marked increase in bias scores for 2020 across all models, with transformer models (both soft and sparse attention) showing particularly pronounced increases, approaching the ideal score of . All of these increases are statistically significant. When a moderate amount of bias () is injected into the training data, the models display a similar pattern: a significant increase in bias scores from 2018 to 2020. In contrast, a significant decrease in bias scores emerges between 2020 and 2024. However, we do not observe a significant change in bias scores between 2018 and 2024, or when the Wikipedia dump is supplemented with synthetic bias. These findings are visually illustrated in Figure 12. We conduct further experiments on raw training data (excluding model effects) that reveal significant bias variations between pre- and post-COVID corpora, as detailed in Appendix C.